A pair of diece is rolled 50 times. Find the propability of odtaining double six at least 3 times.
A-Six appear, P(A)=2/12= 1/6=p ;; q=10/12=5/6
n-50trials , k-3 six

P50(3)=(50)(1/6)^3(5/6)^47 ~~0.01723
................3

Answer by stanbon(75887)

A pair of diece is rolled 50 times. Find the propability of odtaining double six at least 3 times.
A-Six appear, P(A)=2/12= 1/6=p ;; q=10/12=5/6
n-50trials , k-3 six
P50(3)=(50C3)(1/6)^3(5/6)^47 ~~0.01723
==========================================
Binomial with n=50 and p = 1/36
---
P(3<= x <=50) = 1 - binomcdf(50,1/36,2) = 0.1617
=========================
Cheers,
Stan H.